A metal rod with a diameter of 2.30 cm and length of 1.10 m has one end immersed in ice at 32oF and the other end in water at 212oF. If the ice melts at a rate of 1.32 g every 175 s, what is the thermal conductivity of this metal? Identify the metal. Assume there is no heat lost to the surrounding air.
Answer #1

heat conducted =KA*dT/dx

m*L/t =KAdT/dx

1.32 *334/175=K*(pi/4*0.023^2) *(212-32)/1.1

K=37.61W/mK